OpenStudy (anonymous):

Which number is divisible by 4? A. 38 B. 74 C. 116 D. 134

OpenStudy (anonymous):

4/38 4/74 4/116 4/134 do these

OpenStudy (anonymous):

and which ever is a hole number is correct

OpenStudy (anonymous):

Ok thank you!

OpenStudy (anonymous):

none of them are whole numbers tho?

OpenStudy (jadeishere):

Well, technically they're all divisible by 4, but only one of them has a whole-number quotient. Look at 116 again, you can break it down into two parts; 100 + 16 100/4 = 25 16/4 = 4 25 + 4 = 29 so 116 :)

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

need any more help?